# GP-PCS: One-shot Feature-Preserving Point Cloud Simplification with Gaussian Processes on Riemannian Manifolds

### v3

@@ 5.2 Point cloud registration
Table 2: Inlier RMSE and time taken for global and ICP registration. Best results are boldfaced, whilst second-best are underlined.
@@ 6 Conclusion
In this work we have presented a novel, one-shot point cloud simplification algorithm capable of preserving both the salient features and the overall structure of the original point cloud. We reduce the cloud size by up to three orders of magnitude without the need for computationally intensive training on huge datasets. This is achieved via a greedy algorithm which iteratively selects points based on a selection criterion determined by modeling the surface variation over the original point cloud using Gaussian processes with kernels which operate on Riemannian manifolds. We show that our technique achieves competitive results and runtimes when compared to a number of relevant methods, outperforming all baselines tested in terms of mean Hausdorff distance on Lucy, the largest and most complex point cloud we consider, consisting of 14 million points. Our method can also be used to improve the computational efficiency of downstream tasks such as point cloud registration with no negative effects on the empirical performance.
Future work: Whilst Hausdorff distance is a useful metric, it is not the ideal candidate for assessing the feature sensitivity of a simplification algorithm, as it tends to return lower errors for more evenly distributed clouds. Whilst out of the scope of this work, there is a clear need for a well-defined and widely adopted error metric for curvature-sensitive simplification. Currently, the best way to evaluate this is a qualitative visual inspection of the resulting point cloud (or reconstructed mesh). This view is supported by the fact that some recent works employ user studies to evaluate their feature-preserving approaches [28].
In this work we study the setting where we enforce the restriction that the simplified cloud be a subset of the original; as discussed in Section 3.3, a greedy inference scheme is appropriate in this setting. However, this assumption could be relaxed and sparse GPs can be used to perform continuous optimization of the inducing points across the point cloud [15]. This would allow occluded as well as extremely noisy point clouds, where the original observations do not necessarily lie on the true surface of the manifold, to be denoised and/or simplified.
7 Acknowledgements
All the authors would like to thank the University of Antwerp and the University of Manchester for funding this research.
@@ 8 Additional experimental details
Our approach practically has no user-specified parameters, as it generates a simplified cloud based on the desired simplification ratio ($\alpha$). However, we supply to our algorithm (Algorithm 1 in main text) some fixed default values which work well for most point clouds. We set $k_{\text{opt}}=200$ for all experiments as even with a small subset of the original point cloud, the GP typically converges on an optimal set of hyperparameters within 100 iterations. $k_{\text{init}}$ was chosen to be $1/3$ of the target number of points in the simplified cloud, a ratio which empirically works well across all point clouds tested. Finally, $k_{\text{add}}$ is determined adaptively based on $k_{\text{init}}$, $N$ and $M$. Our algorithm is implemented in the PyTorch framework, and whilst the runtimes reported in Table 1 of main manuscript were achieved with GPU acceleration using an NVIDIA A100 with 80GB of RAM, our algorithm can also be run purely on a CPU. All baselines were run on an Intel i7-11800H CPU with 32GB RAM. Our code is available here.
9 Additional visualisations
@@
On the following pages we present additional visualisations to accompany the results presented in the main text. Figure 4 shows the original meshes provided here for a better visual comparison. Figure 5 shows the surface reconstruction results on Lucy. Figure 6 shows the simplified clouds and reconstructed meshes for all techniques on the Armadillo cloud. Figure 7 is a qualitative comparison of the HC and GP-based approaches to performing point cloud registration on the Dragon cloud.
@@
Furthermore, we validate our technique’s feature-sensitive approach on real-world scanning datasets captured using different acquisition devices. Firstly, we use a desk scene point cloud from the NYU Depth V2 dataset, derived from RGBD data acquired using RGB and Depth cameras from Microsoft Kinect. This cloud and the resulting simplification are shown in Figure 8. Secondly, we acquired point clouds from three real-life objects: an angel figure, a human and a bike frame. They were captured using FARO’s scan-in-a-box system, photogrammetry and AliceVision’s Meshroom and Artec 3D’s portable 3D scanner respectively. The three of them when simplified using our method with $\alpha=0.04$, $0.05$ and $0.6$ respectively give results which are shown in Figure 9.
@@
Finally, in Figure 10 we show an extensive comparison of our method with all the mentioned simplification techniques for a range of simplification ratios for the angel point cloud. Again, our method works best for striking a trade-off between enhancing salient features and keeping the low curvature region well-sampled, and hence leads to the best reconstruction results ($\alpha=0.05$) when compared to the original angel mesh (Figure 4). The reconstructions obtained from WLOP, PC-Simp and AIVS simplified clouds are less detailed around salient features (compare the hair, fingers, face, wings and feet) while the one from HC-simplified cloud is quite poor around the torso region of the angel because of less density of points around low curvature regions in the simplified cloud.
@@
Figure 8: GP-based simplification applied on a point cloud derived from real-life NYU Depth V2 Dataset’s desk scene with simplification ratio $\alpha=0.03$.

### v4

@@ 5.2 Point cloud registration
Table 2: Inlier RMSE and time taken for global and ICP registration. Best results are in red, whilst second-best are in green.
@@ 6 Conclusion
In this work we have presented a novel, one-shot point cloud simplification algorithm capable of preserving both the salient features and the overall structure of the original point cloud. We reduce the cloud size by up to three orders of magnitude without the need for computationally intensive training on huge datasets. This is achieved via a greedy algorithm which iteratively selects points based on a selection criterion determined by modeling the surface variation over the original point cloud using Gaussian processes with kernels which operate on Riemannian manifolds. We show that our technique achieves competitive results and runtimes when compared to a number of relevant methods, outperforming all baselines tested in terms of mean Hausdorff distance on Lucy, the largest and most complex point cloud we consider, consisting of approximately 14 million points. Our method can also be used to improve the computational efficiency of downstream tasks such as point cloud registration with no negative effects on the empirical performance.
6.0.1 Future work:
Whilst Hausdorff distance is a useful metric, it is not the ideal candidate for assessing the feature sensitivity of a simplification algorithm, as it tends to return lower errors for more evenly distributed clouds. Whilst out of the scope of this work, there is a clear need for a well-defined and widely adopted error metric for curvature-sensitive simplification. Currently, the best way to evaluate this is a qualitative visual inspection of the resulting point cloud (or reconstructed mesh). This view is supported by the fact that some recent works employ user studies to evaluate their feature-preserving approaches [28].
In this work we study the setting where we enforce the restriction that the simplified cloud be a subset of the original; as discussed in Section 3.3, a greedy inference scheme is appropriate in this setting. However, this assumption could be relaxed and sparse GPs can be used to perform continuous optimization of the inducing points across the point cloud [15]. This would allow occluded as well as outlier-ridden extremely noisy point clouds, where the original observations do not necessarily lie on the true surface of the manifold, to be denoised and/or simplified.
@@ 1 Additional experimental details
Our approach practically has no user-specified parameters, as it generates a simplified cloud based on the desired simplification ratio ($\alpha$). However, we supply to our algorithm (Algorithm 1 in main text) some fixed default values which work well for most point clouds. We set $k_{\text{opt}}=200$ for all experiments as even with a small subset of the original point cloud, the GP typically converges on an optimal set of hyperparameters within 100 iterations. $k_{\text{init}}$ was chosen to be $1/3$ of the target number of points in the simplified cloud, a ratio which empirically works well across all point clouds tested. Finally, $k_{\text{add}}$ is determined adaptively based on $k_{\text{init}}$, $N$ and $M$. Our algorithm is implemented in the PyTorch framework, and whilst the runtimes reported in Table 1 of main manuscript were achieved with GPU acceleration using an NVIDIA A100 with 80GB of RAM, our algorithm can also be run purely on a CPU. All baselines other than Potamias et al. and APES were run on an Intel i7-11800H CPU with 32GB RAM.
2 Additional visualisations and experiments
@@
On the following pages we present additional visualisations and quantitative comparisons to accompany the results presented in the main text. Figure 1 shows the some original meshes provided here for a better visual comparison. Figure 2 shows the surface reconstruction results on Lucy. Figure 3 shows the simplified clouds and reconstructed meshes for all techniques on the Armadillo cloud. Figure 4 is a qualitative comparison of the HC and GP-based approaches to performing point cloud registration on the Dragon cloud.
@@
Furthermore, we validate our technique’s feature-sensitive approach on real-world scanning datasets captured using different acquisition devices. Firstly, we use a desk scene point cloud from the NYU Depth V2 dataset, derived from RGBD data acquired using RGB and Depth cameras from Microsoft Kinect. This cloud and the resulting simplification results from proposed and Potamias et al. methods are shown in Figure 5. Secondly, we acquired point clouds from three real-life objects: an angel figure, a human and a bike frame. They were captured using FARO’s scan-in-a-box system, photogrammetry and AliceVision’s Meshroom and Artec 3D’s portable 3D scanner respectively. The three of them when simplified using our method with $\alpha=0.04$, $0.05$ and $0.6$ respectively give results which are shown in Figure 6. In line with the quantitative results provided in the main text, we also report the mean and maximum Hausdorff distances along with the mean surface variation values of all the aforementioned simplified point clouds in Table 1. As per the given evaluation metrics, this table yet again confirms the fact that our algorithm mostly stays within the top three simplification methods. Yet again, as mentioned in our conclusion, complete disagreement of quantitative and qualitative results for the NYU scene using Potamias et al. and proposed method justifies our argument that visual inspection is the best way to evaluate the feature sensitivity of a method.
@@
Finally, in Figure 7 we show an extensive comparison of our method with all the mentioned simplification techniques for a range of simplification ratios for the angel point cloud. Again, our method works best for striking a trade-off between enhancing salient features and keeping the low curvature region well-sampled, and hence leads to the best reconstruction results ($\alpha=0.05$) when compared to the original angel mesh (Figure 1). The reconstructions obtained from WLOP, PC-Simp, AIVS and Potamias et al. simplified clouds are less detailed around salient features (compare the hair, fingers, face, wings and feet with the original angel mesh in Figure 1) while the one from HC-simplified cloud is quite poor around the torso region of the angel because of less density of points around low curvature regions in the simplified cloud.
2.0.1 Ablation and noise studies:
We conduct experiments to understand the significance of some parameters within our algorithm. These parameters were $k_{init}$, $i$, $k_{opt}$, and $k$ (the number of initial points chosen using FPS, the number of times the hyperparameters of the GP are optimized, the number of points used for hyperparameter optimization, and the number of neighbourhood points used for surface variation estimation). More specifically, while keeping all other parameters fixed and varying just one, we compute the Chamfer distance between the original and the simplified clouds of the Stanford bunny for the simplification ratio ($\alpha$) of 0.2. Moreover, we also quantify and extend the results of simplifying a noisy Armadillo in the main text by providing Chamfer distance between the original Armadillo with varying noise levels (Gaussian noise of standard deviation $n\times d$, where $d$ is the bounding box diagonal length, is added to every point position) and simplified versions of them for the simplification ratio ($\alpha$) of 0.04. All of these results are given in Table 2.
@@ 2.0.1 Ablation and noise studies:
Figure 5: GP-based and Potamias et al. simplification applied on a point cloud derived from real-life NYU Depth V2 Dataset’s desk scene with simplification ratio $\alpha=0.03$.
@@
Table 1: Empirical results for all tested simplification methods for the remaining point clouds. We report the maximum and mean Hausdorff distances between the original meshes, and the meshes reconstructed from the simplified point clouds. Also reported is the average surface variation over each simplified point cloud. Best, second-best and third-best results are in red, green and blue respectively.
Mean Hausdorff Distance ($\downarrow$) Max. Hausdorff Distance ($\downarrow$) Mean Surface Variation ($\uparrow$)
Angel Bike frame Human NYU Angel Bike frame Human NYU Angel Bike frame Human NYU
GP (ours) 0.000172 0.001921 0.196125 0.045540 0.001701 0.019782 4.287749 1.704832 0.0445 0.1021 0.0309 0.0487
HC 0.000270 0.001959 0.395263 0.048864 0.006358 0.019743 10.380267 1.483680 0.0509 0.1003 0.0407 0.0648
WLOP 0.000098 0.006128 0.153166 0.046066 0.000920 0.086360 4.410541 1.671164 0.0294 0.0737 0.0169 0.0403
PC-Simp 0.000150 0.001927 0.369116 0.044758 0.048970 0.020815 12.259072 1.741011 0.0351 0.1031 0.0260 0.0463
AIVS 0.000502 0.004217 0.518831 0.080380 0.002092 0.016377 9.786209 1.754694 0.0321 0.0760 0.0142 0.0530
Potamias et al. 0.000240 0.002117 0.327904 0.012680 0.002247 0.028753 6.091248 0.669750 0.0538 0.1174 0.0376 0.0606
Table 2: Ablation and noise studies on the Stanford bunny and Armadillo respectively. Here, $k_{init}$, $i$, $k_{opt}$, $k$, and $n$ are the number of initial points chosen using FPS, the number of times the hyperparameters of the GP are optimized, the number of points used for hyperparameter optimization, the number of neighbourhood points used for surface variation estimation, and the noise factors respectively (Section 2.0.1).
Ablation studies Noise studies
$k_{init}$ CD $i$ CD $k_{opt}$ CD $k$ CD $n$ CD
500 7.4345e-06 50 3.1976e-06 50 3.2453e-06 5 2.7721e-06 0.0001 1.9456
1000 5.0463e-06 100 3.2049e-06 100 3.2237e-06 10 3.0523e-06 0.0005 1.9532
1500 3.8995e-06 150 3.2471e-06 150 3.2156e-06 15 3.1314e-06 0.001 1.9148
2000 3.2469e-06 200 3.2153e-06 200 3.2529e-06 20 3.2396e-06 0.005 2.1729
3000 2.4893e-06 300 3.2253e-06 300 3.2139e-06 30 3.3033e-06 0.01 3.1361
5000 1.8498e-06 500 3.1996e-06 500 3.2236e-06 50 3.4406e-06 0.02 4.8644